Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day02


Crew 100B Log Book for February 27 2011

Photos of the day





Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent, compared to yesterday

Experiment Status: All the medical, geological, rover, and food experiments began today with interesting results.

The meals today were especially appetizing, since it was a cooking day.

Productivity: We were very productive, since our rover engineer and crew scientist began to really get their experiments started.

EVA’s: We completed two successful EVA’s and training using the ATV’s. Many lessons were learned, that included: how to drive ATV’s, using proper protocols for collecting soil samples, and proper use of the EVA suit backpack, as well as perfecting our communication skills while on EVA’s.

Plan for tomorrow: Begin with doing the medical experiment again on all crew members. Observations of the Sun via the radio telescope will also begin. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will occur. Soil samples will again be collected during two EVA’s tomorrow. One will be on foot, and one will use the ATV’s. Tomorrow will be a non-cooking day


Commander's Report

Balwant Rai Reporting

Today we started with a delicious breakfast of cereal and coffee, planned by our Executive Officer. It was very enjoyable and delicious. Afterwards, we began medical experiments by taking saliva samples via special sensors from all crew members to determine their current stress level. Saliva samples were also taken by another means (via Versi Sal 1 from Oasis Diagnostics). By using the Zephyr Biohraness, vital parameters of active crew members were monitored before, during, and after EVA. As well as pre-sleep and post-sleep. Heart rate variability was also monitored for all crew members. Every crew member also enjoyed using the cognitive activity software.


Today, was a great day for the crew to explore the Martian landscape and practice using the ATV’s. Afterwards, we had a great lunch to follow by Jasdeep and Crystal. This was followed by two EVA’s that involved collecting soil samples for analysis of microbiological activity, as well as extremophiles. Our crew scientist, Crystal, completed all prep work to start observations of the sun, via the radio telescope tomorrow.


During the late afternoon, the cadets worked on getting the rover working and examining how well it maneuvers when being controlled on the outside of the Hab. Our Hab engineer has also been doing a great job at keeping up with the problem pertaining to the water pump. Our chefs are currently working on a good cooked meal tonight, since it is a cooking-day. Jasdeep has also been participating in a unique food study, where she alone is eating a vegetarian diet.


Tomorrow, we plan to begin with doing the medical experiment again on all crew members. The plan is to do this every day for the entire time we are here on Mars. Observations of the Sun via the radio telescope will also begin tomorrow. The rover engineer, Rachel, is planning on continuing her study of the rover and potentially using the drone, weather pending. Mattheiu and Quentin will also join Rachel in examining the feasibility of the rover, here on Mars. Jasdeep will continue her collection of soil samples, with two EVA’s tomorrow. One will be on foot, and one will use the ATV’s. Tomorrow will be a non-cooking day, so it should be a fun filled day here on Mars.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.74

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.42

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): Empty

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 11

· Trailer Potable Water Tank Level (inches from bottom): 2/3 Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading: 424946


Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: Green water

· Condition of plants in Tank 2: Green things

· Condition of plants in Tank 3: Green things

· Greenhab notes/ comments: Water level in Tank 2 and Tank 3 dicreases because Tank 1 isn’t supplied in water.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: no

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 65

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: We didn’t succeed in connecting the printer to our computer.


General Engineering

1. Notes/Comments:

Clear


Science Report

Rachel Dompnier Reporting

Today I have made tests to prepare myself for using the rover and the drone.


First, with the help of Quentin Bourgès, I familiarized myself with operating the rover. Quentin was on an EVA outside the HAB, where he guided me and made sure everything worked in case of a human error by me. I operated the rover from the green HAB where I was first able to easily see the rover outside by opening the door and then later operating the rover from inside with the camera and the computer.



I confronted some difficulties with controlling the rover. This occurred when I was pushing the joystick completely forward, and instead of moving steadily forward the rover jerked forward. I do not know what the problem may be but it could be due to a connection problem, low power, or it may not be a problem at all. Another problem that I encountered was a delay in the recording of the rover’s cameras. I found it difficult to avoid obstacles with this delay. Also I had a hard time determining at what distance the rover was from obstacles, which resulted in a lack of maneuverability.


After this test, we made sure we downloaded the correct application for the iPad that was used to control the drone. We found that we did have the correct application. Because of the wind we were not able to go outside with the drone, therefore we learned how to pilot the drone on the ground floor of the Hab. There were no problems operating the drone today.


Science Report (Biology)

Jasdeep Kaur Reporting

Time: 0230-sample collection

Crew Members:

Science Team: 3 crew members,

EVA mission supporters : Two crew Members

Site Location: Hab

Sample Collection Site: N 38.40717, W 110.78620

Elevation: 1362 meters

Transit Mode: ATV

Objective: To find out the microbial and extremophile activity

Sample Collection Tools:

Sterile 50mL centrifuge tubes/vials with screw cap

Alcohol (96%)

Lighter (to flame alcohol)

Nitrile gloves

Sterile stainless steel trowel or spatula

Sterile plastic bag or sterile glass (250ml/500ml)

Cooler with blue ice or freezer for storing samples

Sharpie permanent marker

Small sealable plastic sandwich bags for storing centrifuge tubes after sampling

GPS

Field sheet, pencil/pen


Sampling Depth: In general, soil samples were taken usually 0 to 1 inches in depth. A constant depth was maintained for all samples collected for best results.


Search for sample collection: A specific site was searched by three crew members who went for an EVA training in morning at 11.30am. A particular area with greenish area was suspected of growth of bacterial and fungal colonies. This was recommended for sample collection by biologist Jasdeep Kaur on ATV mission.


Site-Specific Collection: Site-specific soil sampling was performed. The objective was to allocate the specific areas where more findings of extremophiles and bacterial activity were suspected. A field was mapped in a specific direction from hub and divided into blocks or cells to form a grid. Using GPS software, specific green coloured location within each grid cell was identified and sampled. Samples were placed in in a clearly-labeled Soil Sample Bag and preserved under refrigeration.


Recording findings: Another crew member Crystal Lantham recorded all specific finding from GPS and noted down using a pen and paper.


Method of collection: Special sterile gloves were used to collect sample. A stainless steel spoon was cleaned and then sterilization was carried out with alcohol. Special centrifuged tubes were used to collect sample and pictures of sample collection were taken. These tubes were sealed and brought back to Hab which were centrifuged thereafter for preparing homogeneous mixture for analysis. Samples were preserved in refrigeration.


Preservation: The preservation of samples is best carried out in cold temperature; this further preserves any further contamination. All samples were preserved in refrigerator of hab on first level.


Science Report (Medical)

Balwant Rai reporting

Physiological Variables of EVA Participants of Mars Analogue Mission


Placing a human into free space, outside the protective confines of the space vehicle, has become an important component of space travel and exploration. An Extravehicular Activity (EVA) is a protocol that requires the astronaut to leave the aircraft to perform various duties such as repairs, observations, and experiments. One of the most important aspects of preparing for an EVA is to determine the physical requirements of the participants. Although the astronaut corps consists of individuals with specific skills, not everyone may be suitable for EVA activity. It has been reported that extravehicular activity (EVA) exposes the crew member to substantial physical stress. Use of EVA suits during Mars analogue work, particularly those with helmets exhibiting poor ventilation, has been reported to cause increased exertion, expressed as heart rate response, during session of physical work. It is of utmost importance that the crew of a craft that is to land on Mars be physically fit in order to handle their work tasks and to be prepared for possible emergency situations. A journey to Mars will take months to complete and as such an appropriate exercise regime must be in place to ensure the Safety and productivity of the crew. To the best my knowledge no evaluation made of the physical work involved in Mars analogue EVA activity. Hence this study is planned the effects of EVA on human physiology. All crews members were selected for this study. The Heart Rate, breathing rate and VO2 max were measured before and after EVA by Zephyr BioHarness using OmniSense live and OmniSense analysis. Heart rate, breathing rate, and VO2 levels were significantly increased. So Marsonaut should be required to participate in physical activity and making a proper space suit that will prevent these adverse effects.


Science Report (Dietitian)

Jasdeep Kaur reporting

A special science experiment on diet pattern was conducted. The main objective behind this was to refer to findings by "ten percent law," also often called the "law of tens" refers to the design that at each transfer of energy through a trophic structure, only a small percent of the energy remains available for use by the organism in the next level up in the system. The majority of the energy is lost to the surroundings in forms such as waste heat. Trophic structure refers to what is often called the food chain or food web, wherein green plants (the producer level) capture energy from the sun through the process of photosynthesis and transform it into energy available to other living organisms (the consumer levels). The ten percent law suggests or implies that exactly 90% of the energy is lost in the transfer at each trophic level, and that only 10% is passed on as usable biological energy.


SUN (1000 joules of light energy)


PLANTS >>>>>>>>> ANIMALS (fed on plants) >>>>>>>> ANIMALS (fed on animals)


(10 joules as food) >>>>>>> (1 joule energy available as food) >>>>>>> (0.1 joule energy available as food)


This experiment refers to restrict one of crew member on vegetarian diet and others mixed diet to evaluate the activity of work. For vegetarian diet equal proportion of grams and vegetables were served to balance the calorie intake. Fruits were also included. Other crew members were served with mixed diet consisting of vegetable soups and salami sandwiches. Activity test will be performed on 3rd day to figure out the changes in crew members. This test will include 2 subjects who had vegetarian and mixed diet to find out the differences. Further-on, one crew member will be restricted to non-veg diet, one vegetarian diet and third one on mixed diet and activity test will be performed on 6th day.


EVA 2 Report

Crystal Latham Reporting

EVA #2, 14:30-15:00

Crew Members: Matthieu Ansart, Crystal Latham, Jasdeep Kaur

Site Location: N 38.40717, W 110.78620

Transit Mode: ATVs

Objectives: GeoMission 1: to collect rock/soil samples to look for evidence of microbial and extremophile life.

Results: Collected a sample of greenish-white rock/ and soil that was in an area that exhibited characteristics of water flowing, as well as ground fractures in an area that had a very weird color.

Lessons Learned: Be sure to follow all protocols when collecting samples.


EVA 3 Report

Crystal Latham Reporting

EVA #3, 16:00- 16:45

Crew Members: Balwant Rai, Rachel Dompnier , Quentin Bourges

Site Location: N 38.40886, W 110.14756     

Transit Mode: ATVs

Objectives: GeoMission 1: to collect rock/soil samples to look for evidence of microbial and extremophile life.

Results: Collected a sample of white soil that was in an area that exhibited characteristics of water flowing, as well as ground fractures.

Lessons Learned: There is potential to drive the ATV with only the space helmet, if all protocols are followed.


Chef's Report

Quentin Bourges Reporting

1a) Today was

- A cooking day

- Cooking: Chrystal Latham and Jasdeep Kaur

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

1 package of Re-hydrated chicken and pastas (16oz)

Leftover from lunch (Santa Fe Chipotle plus mixed pepper, broccoli and corn)

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Nothing

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

No

5) Lessons learned (if any)

It may be obvious but make sure every body is ready to eat when diner’s on the table so it doesn’t get cold.

6) Comments/questions for Kim and Jean, the food study investigators

No


Journalist Report

Crystal Latham Reporting

It is crew 100B’s first full day on Mars and it has been action packed. We started the day off with some medical tests that included taking saliva samples from crew members and monitoring their vitals with some very advanced equipment and software.

A few members of the crew were not very comfortable with riding the ATV’s yesterday, so today we had a practice session. It was a very successful session that really allowed all the crew members to explore the surrounding Martian landscape. The rest of the day consisted of two EVA’s, using the ATV’s, to collect soil samples that will be tested for evidence of extremophiles and microbiological activity.

Today was very busy, and it showed in the results from one of the medical experiments that showed some of us are quite stressed. Today is a cooking day, where our Martian gourmet chefs are Jasdeep and I. We are really looking forward to a nice meal to ease our stress from the day.

Tonight the plan is to begin calibration of the radio telescope, and tomorrow beginning our first observations of the radio Sun. It has been a great day on Mars.